Suppose we do not know the population standard deviation. What is the margin of error when the standard error is 4, and the confidence coefficient is .98 and we are working with a sample of 35 observations?
Answer #1

here for (0.98 confidence interval and (n-1=34 )degree of freedom; critial t =2.441

therefore margin of error =t*Std error =4*2.441=9.764